Opened 11 years ago

Last modified 7 years ago

#175 assigned flight processing

GB08/08, flight day 176b/2008, Wytham_Woods

Reported by: mggr Owned by: adbe
Priority: alpha 4 high Milestone: Reprocessing
Component: Processing: general Keywords:
Cc: jatt Other processors:

Description (last modified by mggr)

Data location: ~arsf/arsf_data/in_progress/2008/flight_data/uk/GB08_08-2008_176b_Wytham_Woods

Data arrived from ARSF via SATA disk N, 2/July/2008.

Scientific details: see ~/arsf/arsf_data/in_progress/2008/ARSF_Applications/GB_and_NL/GB08-08 Science Case.pdf

PI: R Harding

Notes:

  • originally delivered as GB08/05, but changed project code per discussion #151.
  • only applanix required for this project, deliver raw data and ancillary files to Andrew Wilson.
  • eagle data missing from initial delivery

Sensors:

  • ATM
  • Eagle (missing in initial delivery, arrived 7/July)
  • Hawk

Quicklook

TBD.

Change History (10)

comment:1 Changed 11 years ago by mggr

  • Description modified (diff)

In a phone call with Gary, he mentioned that there was a problem with the Eagle data (disk corruption on the original disk?). Phil was going to try a recovery, which Gary said normally works.

Suggest delivering without Eagle data for now (applanix only processing anyway) and following up when it's available.

comment:2 Changed 11 years ago by mggr

  • Owner set to jatt

comment:3 Changed 11 years ago by mggr

  • Description modified (diff)

Updated PI per #151 comments

comment:4 Changed 11 years ago by benj

  • Cc jatt added

Processed by jatt, rsynced to ~arsf/arsf_data/in_progress/2008/flight_data/uk/GB08_08-2008_176b_Wytham_Woods

comment:5 Changed 11 years ago by mggr

  • Description modified (diff)

Eagle arrived 7/July, got onto disk prior to delivery.

comment:6 Changed 9 years ago by benj

  • Resolution set to fixed
  • Status changed from new to closed

NEODC archived 31/07/2009, closing ticket.

comment:7 Changed 7 years ago by benj

  • Resolution fixed deleted
  • Status changed from closed to reopened

Level 1 data doesn't seem to have gotten from AKW to NEODC. Eagle/Hawk need to be reprocessed and sent to NEODC, please notify Charles George (ctg@…) when done.

comment:8 Changed 7 years ago by benj

  • Owner changed from jatt to adbe
  • Status changed from reopened to assigned

comment:9 Changed 7 years ago by adbe

Created Delivery

comment:10 Changed 7 years ago by adbe

Ready to deliver at:

/users/rsg/arsf/workspace/GB08_08-2008_176b_Wytham_Woods/processing/delivery/GB08_08-176b-hyperspectral-20120201

Note: See TracTickets for help on using tickets.